OpenStudy (anonymous):

Limit as x approaches 1 of: (sqrt(x) - 1)/(sqrt(x-1)) I can't use L'Hôpital's rule because we haven't learned it yet.

zepdrix (zepdrix):

\[\huge \lim_{x \rightarrow 1}\frac{\sqrt x - 1}{\sqrt{x-1}}\]Multiplying top and bottom by the CONJUGATE of the top, gives us: \[\huge \lim_{x \rightarrow 1}\frac{\sqrt x - 1}{\sqrt{x-1}}\left(\frac{\sqrt x + 1}{\sqrt x + 1}\right)\]

zepdrix (zepdrix):

Remember what happens when you multiply conjugates? :D You get the difference of squares.

zepdrix (zepdrix):

Hmmmm I'll bet this puts us on the right track :o hopefully..

zepdrix (zepdrix):

Maybe not.. hmmmmm

OpenStudy (anonymous):

Well, the denominator is still zero after 1 is plugged in.

zepdrix (zepdrix):

@Zarkon to the rescue? :D hehe

OpenStudy (zarkon):

\[(\sqrt{x}-1)(\sqrt{x}+1)=x-1\] \[\frac{x-1}{\sqrt{x-1}}=\sqrt{x-1}\]

OpenStudy (zarkon):

\[\frac{\sqrt{x-1}}{\sqrt{x}+1}\to0\text{ as }x\to1\]

OpenStudy (anonymous):

err... Why?

OpenStudy (zarkon):

\[\lim_{x \rightarrow 1}\frac{\sqrt x - 1}{\sqrt{x-1}}\left(\frac{\sqrt x + 1}{\sqrt x + 1}\right)\] \[=\lim_{x \rightarrow 1}\frac{x-1}{\sqrt{x-1}(\sqrt x + 1)}\] \[=\lim_{x \rightarrow 1}\frac{\sqrt{x-1}}{\sqrt x + 1}=\frac{0}{2}=0\]

OpenStudy (anonymous):

wow, I'm blind. This is pretty obvious lol. Thanks for help!